OpenStudy (anonymous):

What is the magnitude and direction of Vector GH if G(2, -2) and H(7, 6)?

OpenStudy (anonymous):

magnitude: 9.43 units; direction 32.01 magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99 magnitude: 6.4 units: direction: 32.01

OpenStudy (jdoe0001):

is that G+H? or \(\bf G \cdot H\)?

OpenStudy (anonymous):

its a vector

OpenStudy (anonymous):

so plus i think

OpenStudy (jdoe0001):

well, the G+H will be = <2+7, -2+6> => <9, 4>

OpenStudy (jdoe0001):

\(\bf \text{magnitude of a vector} \\ ||v|| = <a, b> = \sqrt{a^2+b^2}\)

OpenStudy (jdoe0001):

and the angle or direction is going towards comes from the angle from \(\bf \cfrac{v}{||v||}\)

OpenStudy (jdoe0001):

well, that wont' give you much of an angle, but if you take the tangent of the 2 components, should

OpenStudy (anonymous):

ok thanks

OpenStudy (jdoe0001):

thus the GH = <9, 4> is moving at the angle of $$ tan(\theta)= \frac{4}{9}\\ \theta = tan^{-1}\pmatrix{\frac{4}{9}} \implies 23.96^o $$

OpenStudy (anonymous):

right...